calculate the volume of oxygen needed to completely react with 150ml of carbon monoxide. assume al the volumes are measured at same temperatures and pressures.


1
Expert's answer
2021-11-02T02:07:05-0400

2CO + O2 = 2CO2

According to the reaction, n (O2) = 1/2 x n (CO).

n = V / 22.4


n (CO) = 0.15 / 22.4 = 0.007 mol

n (O2) = 0.007 / 2 = 0.0035 mol


V (O2) = 0.0035 x 22.4 = 0.078 L = 78 ml